OpenStudy (anonymous):

What is the simplified form of 2/x^2-x minus 1/x

OpenStudy (nirmalnema):

x^2-x-1/x is whole in the denominator

ganeshie8 (ganeshie8):

\(\large \frac{2}{x^2-x} - \frac{1}{x} \)

ganeshie8 (ganeshie8):

\(\large \frac{2}{x(x-1)} - \frac{1}{x} \)

OpenStudy (anonymous):

then times the other side by x-1 ?

ganeshie8 (ganeshie8):

Yes !!!

ganeshie8 (ganeshie8):

\(\large \frac{2}{x(x-1)} - \frac{x-1}{x(x-1)}\)

ganeshie8 (ganeshie8):

since denominators are common, we can add numerators

ganeshie8 (ganeshie8):

\(\large \frac{2-(x-1)}{x(x-1)} \)

ganeshie8 (ganeshie8):

\(\large \frac{2-x+1}{x(x-1)} \)

ganeshie8 (ganeshie8):

\(\large \frac{3-x}{x(x-1)} \)

OpenStudy (nirmalnema):

3-x/x(x-1)

OpenStudy (anonymous):

your an amazing help ! i got a 100% thank you sooo muchh ! (:;

ganeshie8 (ganeshie8):

wow ! thats brilliant !! congratulations xD

ganeshie8 (ganeshie8):

you should partyy :D

OpenStudy (anonymous):

Lol (: & Your invitedd ! @ganeshie8 i have to run an arron. itll take like 20 min . but when i get back are you willing to help in another assignment ? its olny 8 questions ? (:::: pleaseee.

ganeshie8 (ganeshie8):

ahh ty so much :) sure il be happy to assist the brilliant student like you :)

OpenStudy (anonymous):

Whyy thank you. (: . and ill tagg you as soon as back !

ganeshie8 (ganeshie8):

okie :)